# VoiceSculptor: Your Voice, Designed By You

### v1

@@ 3.1 Evaluation On InstructTTSEval-Zh Benchmark
Our training data predominantly consists of Chinese instruction–speech pairs. Therefore, we focus our evaluation on the model’s instruction-following and controllability performance in Chinese, which allows for a fair and consistent assessment aligned with the training distribution. We emphasize that this choice does not imply any inherent limitation of the proposed method to Chinese. Instead, Chinese is adopted as a representative language to demonstrate the feasibility and effectiveness of our approach, which can be naturally extended to other languages given appropriate instruction data.
To this end, we employ InstructTTSEval-Zh, a Chinese instruction-based TTS evaluation benchmark designed to measure how well a model follows natural-language instructions in speech synthesis. The benchmark evaluates multiple aspects of instruction controllability, including attribute perception and synthesis accuracy (APS), description–speech consistency (DSD), and response precision (RP). These metrics are computed by synthesizing speech from instruction prompts and assessing the generated audio using a unified evaluation protocol with a large language model as the evaluator.
@@
Model APS (%) DSD (%) RP (%) AVG (%)
Gemini 2.5-Flash* 88.2 90.9 77.3 85.4
Gemini 2.5-Pro* 89.0 90.1 75.5 84.8
@@
Under this evaluation setting, VoiceSculptor-VD consistently outperforms all open and instruction-tuned baselines across the majority of metrics. In particular, it achieves the best performance in APS and RP, indicating superior accuracy in attribute perception and instruction-to-acoustic rendering within the voice design module. These results demonstrate that, when combined with retrieval-based instruction grounding, the VD module of VoiceSculptor is highly effective at translating abstract natural language descriptions into fine-grained and controllable vocal characteristics, which constitutes the core objective of voice design.
@@
Although Gemini 2.5-Flash and Gemini 2.5-Pro achieve strong overall scores, they are commercial proprietary models and rely on closed-source infrastructures. In contrast, VoiceSculptor achieves state-of-the-art performance among open-source instruction-following TTS models, while maintaining a unified, instruction-driven voice design framework.
@@
Compared with MiMo-Audio-7B-Instruct, which achieves slightly higher performance on DSD, VoiceSculptor-VD shows clear advantages in APS and RP, leading to the highest overall AVG score among open systems. This indicates that the voice design module of VoiceSculptor prioritizes precise and consistent attribute control over isolated stylistic similarity, resulting in more faithful and reliable instruction-following behavior at the voice design stage. The substantial performance margin over VoxInstruct and GPT-4o-Mini-TTS further highlights the effectiveness of the design choices specific to the VD module, including fine-grained attribute modeling, instruction-aware training, and robust scaling strategies.
@@
To further account for the requirements of downstream speech synthesis tasks, particularly the preservation of stylistic characteristics conveyed by prompt waveforms, we additionally report results for VoiceSculptor-VD and VC. In this setting, the prompt waveform generated by the voice design module, together with the corresponding test text, is fed into the CosyVoice2 model to perform downstream speech synthesis. This evaluation protocol allows us to directly assess whether the vocal style specified and generated at the voice design stage can be faithfully retained when transferred to a subsequent synthesis model.Experimental results indicate that the stylistic attributes encoded in the prompt waveforms produced by VoiceSculptor-VD are largely preserved in the downstream synthesized speech. Despite the introduction of an additional synthesis stage and a different model architecture, the generated speech maintains strong consistency with the intended vocal characteristics, demonstrating effective style transfer and robustness of the designed prompt representations. These findings suggest that the voice design outputs of VoiceSculptor-VD are not only effective in isolation, but also serve as reliable and reusable conditioning signals for downstream text-to-speech systems, thereby supporting practical deployment scenarios where voice design and speech synthesis are decoupled.
@@
Overall, these results demonstrate that the voice design module of VoiceSculptor achieves state-of-the-art performance among instruction-following TTS systems on the voice design benchmark. The consistent improvements across multiple complementary metrics confirm that the VD module can accurately interpret natural language instructions and reliably generate the desired vocal attributes, establishing VoiceSculptor’s voice design component as a strong and practical solution for controllable voice design in instruction-driven text-to-speech systems.
@@ 3.2 Scaling Study On Model Size And Data Size
Evaluating instruction-following performance using large-scale automated benchmarks requires repeated calls to proprietary models such as Gemini, which significantly increases evaluation cost and latency. To enable rapid and iterative validation of scaling trends, we construct an internal lightweight subjective evaluation benchmark consisting of 100 carefully curated test instructions and their corresponding synthesized speech outputs.
@@
Training Configuration IMOS APS (%) DSD (%) RP (%) AVG (%)
1B / $SFT_{Data1}$, 3 epoch 3.09 51.3 48.2 35.7 45.1
3B / $SFT_{Data1}$, 3 epoch 3.24 59.2 53.1 39.4 50.6
1B / $SFT_{Data2}$, 3 epoch 3.35 61.5 55.9 45.9 54.4
3B / $SFT_{Data2}$, 3 epoch 3.58 72.4 60.0 52.8 61.8
3B / $CPT_{Data4}$, 2 epoch and $SFT_{Data3}$, 3 epoch 3.67 75.7 64.7 61.5 67.6
We conduct a human listening study focusing on instruction-following capability, measured by Instruction-following Mean Opinion Score (IMOS). IMOS specifically assesses how well the synthesized speech adheres to the semantic and stylistic requirements expressed in the natural language instruction, rather than overall audio quality alone. A total of 33 human listeners participated in the evaluation. Each listener is randomly assigned 10 audio samples drawn from the 100-sample test set, ensuring that each test sample is evaluated by multiple listeners while avoiding listener fatigue. Listeners are asked to provide subjective scores based on how accurately the synthesized speech follows the given instruction, using a standardized MOS-style rating scale. The final IMOS score is computed by averaging all listener ratings across samples.
The results in 3 demonstrate that model performance consistently benefits from both increased model capacity and enlarged, more diverse training data. Under identical Supervised Fine-Tuning (SFT) settings, scaling the model from 1B to 3B parameters yields clear improvements across all metrics, indicating stronger representation and generalization capabilities. For a fixed model size, expanding SFT data from in-the-wild only to larger mixtures with internal data leads to substantial gains, highlighting the importance of data scale and diversity. Finally, incorporating large-scale emotion-aware continual pre-training (CPT) on $CPT_{Data4}$ prior to SFT achieves the best overall performance, suggesting that CPT provides a more favorable initialization and enables the model to better exploit downstream supervised data for the voice design task.Figure 3 and Figure 4 further corroborate these findings by showing consistently lower validation loss for larger models, and richer training data throughout training.
Overall, these results indicate that incorporating CoT fine-grained attribute tokens, together with stochastic attribute dropout, provides a robust and scalable enhancement, yielding consistent validation loss reduction across different model sizes.
@@ 3.3 Ablation Study On CoT-based Fine-grained Attribute Tokens
The ablation study on CoT-based fine-grained attribute tokens validates the effectiveness of introducing auxiliary attribute tokens into the voice design framework. As shown by the consistent improvements across all evaluation metrics, these tokens enable the model to better capture and utilize fine-grained attribute information, leading to more accurate and controllable speech generation. Beyond quantitative gains, the use of CoT-based attribute tokens facilitates finer-grained control over voice characteristics, allowing more precise manipulation of individual attributes. This enhanced controllability is further demonstrated through qualitative examples on our demo page, where the impact of fine-grained attribute conditioning can be clearly observed.
@@ 3.4 Ablation Study Of Text Cross-Entropy Loss
Based on the results in Table 5, introducing the text-side cross-entropy (CE) loss during training proves to be highly effective. By jointly modeling the instruction text and audio tokens within a unified training objective, the model achieves consistent and significant improvements across all evaluation metrics. Compared to the setting without text CE loss, incorporating text CE loss leads to notable gains in IMOS, APS, DSD, RP, and the overall AVG score, indicating enhanced alignment between textual instructions and generated speech.
@@
Setting IMOS APS (%) DSD (%) RP (%) AVG (%)
VoiceSculptor-VD 3.67 75.7 64.7 61.5 67.6
VoiceSculptor-VD w/o Text CE Loss 3.42 67.9 59.4 58.2 61.8
These results suggest that explicitly supervising the text modality encourages the model to better capture long-range contextual dependencies and semantic intent conveyed by instructions, rather than treating text merely as auxiliary conditioning. As a result, the model develops stronger context understanding and instruction-following capabilities, which directly translate into improved controllability and perceptual quality in the voice design task.
@@ 3.5 Ablation Study Of RAG
Table 6 demonstrates the necessity and effectiveness of incorporating RAG into the proposed VoiceSculptor-VD framework. Enabling RAG leads to substantial improvements across all evaluation metrics, with especially pronounced gains in APS (+7.1%), RP (+13.0%), and the overall AVG score (+8.2%). These results indicate that external retrieval provides critical contextual and attribute-related information that the model alone may fail to infer reliably from the instruction text, thereby significantly enhancing controllability and instruction adherence in the voice design task.
@@
Setting IMOS APS (%) DSD (%) RP (%) AVG (%)
VoiceSculptor-VD 3.67 75.7 64.7 61.5 67.6
VoiceSculptor-VD w/o RAG 3.39 68.6 61.1 48.5 59.4
At the same time, the large performance gap between the RAG and non-RAG settings also reveals inherent limitations of the current model. In particular, the model exhibits relatively limited text understanding and generalization when relying solely on its internal representations, making it sensitive to instruction phrasing and less robust to unseen or complex textual descriptions. By supplementing the model with retrieved examples and structured attribute information, RAG effectively compensates for these shortcomings.
In future work, we plan to address these limitations more fundamentally by strengthening the model’s text understanding capabilities. Specifically, we will explore incorporating large-scale text data during the pre-training stage to better preserve and enhance linguistic representations, performing instruction data augmentation to improve robustness and generalization, and adopting more semantically expressive audio representations by replacing the current xcodec2 with a more semantically expressive codec, thereby alleviating the model’s reliance on external retrieval while maintaining strong instruction-following performance.
@@ 4 Conclusion
In this work, we present VoiceSculptor, a unified instruction-driven voice design framework that enables fine-grained and controllable speech synthesis through natural language. Extensive evaluations on the InstructTTSEval-Zh benchmark demonstrate that VoiceSculptor-VD achieves state-of-the-art performance among open-source instruction-following TTS systems, consistently outperforming strong baselines across multiple complementary metrics. Our scaling study confirms that instruction-following capability benefits predictably from increased model capacity, richer training data, and staged training strategies with large-scale continual pre-training. Through systematic ablation studies, we further validate the effectiveness of key design choices, including CoT-based fine-grained attribute tokens, text-side cross-entropy supervision, and retrieval-augmented generation, each of which contributes to improved instruction understanding, controllability, and robustness. In particular, CoT-based attribute modeling enables more precise and interpretable control over vocal characteristics, while RAG effectively compensates for the model’s limited text generalization by providing semantically aligned in-domain guidance. Finally, we demonstrate that the voice design outputs of VoiceSculptor can be reliably transferred to downstream speech synthesis models, supporting practical deployment scenarios where voice design and speech generation are decoupled. Together, these results establish VoiceSculptor as a scalable and effective solution for instruction-driven voice design.
@@ 5 Ethics Statement
Do not use this model for unauthorized voice cloning, impersonation, fraud, scams, deepfakes, or any illegal or malicious activities. Ensure compliance with local laws and regulations when using this model and uphold ethical standards. The developers assume no liability for any misuse of this model. Important clarification regarding generated voices: As a generative model, the voices produced by this system are synthetic outputs inferred by the model, not recordings of real human voices. The generated voice characteristics do not represent or reproduce any specific real individual, and are not derived from or intended to imitate identifiable persons. We advocate for the responsible development and use of AI and encourage the community to uphold safety and ethical principles in AI research and applications.

### v2

@@ 3.1 Evaluation on InstructTTSEval-Zh Benchmark
Our training data predominantly consists of Chinese instruction–speech pairs. Therefore, we focus our evaluation on the model’s instruction-following and controllability performance in Chinese, enabling a fair and consistent assessment aligned with the training distribution. We emphasize that this choice does not imply any inherent limitation of the proposed method to Chinese. Instead, Chinese is adopted as a representative language to demonstrate the feasibility and effectiveness of our approach, which can be naturally extended to other languages given appropriate instruction data.
To this end, we employ InstructTTSEval-Zh, a Chinese instruction-based TTS evaluation benchmark designed to measure how well a model follows natural-language instructions in speech synthesis. The benchmark evaluates multiple aspects of instruction controllability, including Acoustic-Parameter Specification (APS), Descriptive-Style Directive (DSD), and Role-Play (RP). These metrics are computed by synthesizing speech from instruction prompts and assessing the generated audio using a unified evaluation protocol with a large language model as the evaluator.
@@
[…]
Gemini 2.5-Flash* (Comanici et al., 2025) 88.2 90.9 77.3 85.4
Gemini 2.5-Pro* (Comanici et al., 2025) 89.0 90.1 75.5 84.8
@@
In this evaluation setting, VoiceSculptor-VD consistently surpasses all open-source and instruction-tuned baselines across most evaluation metrics. Notably, it attains the best performance on APS and RP, reflecting its superior proficiency in both attribute perception and role-play within the voice design module. These findings indicate that, when integrated with retrieval-based instruction grounding, the VD module of VoiceSculptor is highly effective at mapping abstract natural language descriptions to precise and controllable vocal characteristics, which aligns with the fundamental goal of voice design.
@@
Although Gemini 2.5-Flash and Gemini 2.5-Pro achieve strong overall scores, they are commercial, proprietary models that rely on closed-source infrastructure. In contrast, VoiceSculptor achieves state-of-the-art (SOTA) performance among open-source instruction-following TTS models, while maintaining a unified, natural-language-driven voice design framework.
Compared with MiMo-Audio-7B-Instruct, which achieves slightly higher performance on DSD, VoiceScul-
@@
ptor-VD shows clear advantages in APS and RP, leading to the highest overall AVG score among open systems. This indicates that the voice design module of VoiceSculptor prioritizes precise and consistent attribute control over isolated stylistic similarity, resulting in more faithful and reliable instruction-following behavior at the voice design stage. The substantial performance margin over VoxInstruct and GPT-4o-Mini-TTS further highlights the effectiveness of the design choices specific to the VD module, including fine-grained attribute modeling, instruction-aware training, and robust scaling strategies.
@@
To further account for the requirements of downstream speech synthesis tasks, particularly the preservation of stylistic characteristics conveyed by prompt waveforms, we additionally report results for VoiceSculptor-VD and VC. In this setting, the prompt waveform generated by the voice design module, along with the corresponding test text, is fed into the CosyVoice2 model for downstream speech synthesis. This evaluation protocol allows us to directly assess whether the style specified and generated at the voice design stage can be faithfully retained when transferred to a subsequent synthesis model. Experimental results indicate that the stylistic attributes encoded in the prompt waveforms produced by VoiceSculptor-VD are preserved mainly in the downstream synthesized speech. Despite the introduction of an additional synthesis stage and a different model architecture, the generated speech maintains strong consistency with the intended characteristics, demonstrating effective style transfer and robustness of the designed prompt representations. These findings suggest that the voice design outputs of VoiceSculptor-VD are not only effective in isolation but also serve as reliable, reusable conditioning signals for downstream text-to-speech systems, thereby supporting practical deployment scenarios in which voice design and speech synthesis are decoupled.
@@
Overall, these results demonstrate that the voice design module of VoiceSculptor achieves state-of-the-art performance among instruction-following TTS systems on the voice design benchmark. The consistent improvements across multiple complementary metrics confirm that the VD module can accurately interpret natural-language instructions and reliably generate the desired vocal attributes, establishing VoiceSculptor’s voice design component as a strong, practical solution for controllable voice design in natural-language-driven text-to-speech systems.
@@ 3.2 Scaling Study on Model Size and Data Size
Evaluating instruction-following performance using large-scale automated benchmarks requires repeated calls to proprietary models such as Gemini, which significantly increases evaluation cost and latency. To enable rapid, iterative validation of scaling trends, we construct an internal, lightweight subjective evaluation benchmark comprising 100 carefully curated test instructions and their corresponding synthesized speech outputs.
@@
[…]
We conduct a human listening study focusing on instruction-following capability, measured by Instructi-
on-following Mean Opinion Score (IMOS). IMOS assesses explicitly how well synthesized speech adheres to the semantic and stylistic requirements expressed in the natural-language instruction, rather than overall audio quality alone. A total of 33 human listeners participated in the evaluation. Each listener is randomly assigned 10 audio samples drawn from the 100-sample test set, ensuring that each test sample is evaluated by multiple listeners while avoiding listener fatigue. Listeners are asked to provide subjective scores based on how accurately the synthesized speech follows the given instruction, using a standardized MOS-style rating scale. The final IMOS score is computed by averaging all listener ratings across samples.
Table 2 further details the duration and source composition of each training dataset, providing a clearer view of how data scale and diversity evolve across different training stages. The results in table 3 demonstrate that model performance consistently benefits from both increased model capacity and enlarged, more diverse training data. Under identical Supervised Fine-Tuning (SFT) settings, scaling the model from 1B to 3B parameters yields clear improvements across all metrics, indicating stronger representation and generalization capabilities. For a fixed model size, expanding SFT data from in-the-wild to larger mixtures with internal data yields substantial gains, highlighting the importance of data scale and diversity. Finally, incorporating large-scale emotion-aware continual pretraining (CPT) on $CPT_{Data4}$ before SFT achieves the best overall performance, suggesting that CPT provides a more favorable initialization and enables the model to better exploit downstream supervised data for the voice design task. Figure 3 and Figure 4 further corroborate these findings by showing consistently lower validation loss for larger models, and richer training data throughout training.
Increasing model capacity and enriching the quality and diversity of training data are fundamental to improving speech instruction-following. Building on this, the CPT stage offers a more robust initialization for the SFT stage, thereby significantly boosting overall performance on the Voice Design task.
@@ 3.3 Ablation Study on CoT-based Fine-grained Attribute Tokens
The ablation study on CoT-based fine-grained attribute tokens validates the effectiveness of introducing auxiliary attribute tokens into the voice design framework. As shown by consistent improvements across all evaluation metrics, these tokens enable the model to capture better and utilize fine-grained attribute information, leading to more accurate and controllable speech generation. Beyond quantitative gains, the use of CoT-based attribute tokens facilitates finer-grained control over voice characteristics, allowing more precise manipulation of individual attributes. This enhanced controllability is further demonstrated through qualitative examples on our demo page, where the impact of fine-grained attribute conditioning is clearly evident.
@@ 3.4 Ablation Study of Text Cross-Entropy Loss
Based on the results in Table 5, introducing the text-side cross-entropy (CE) loss during training is highly effective. By jointly modeling the instruction text and audio tokens within a unified training objective, the model achieves consistent and significant improvements across all evaluation metrics. Compared to the setting without text CE loss, incorporating text CE loss yields notable gains in IMOS, APS, DSD, RP, and the overall AVG score, indicating greater alignment between textual instructions and generated speech.
@@
[…]
These results suggest that explicitly supervising the text modality encourages the model to capture better long-range contextual dependencies and the semantic intent conveyed by instructions, rather than treating text merely as auxiliary conditioning. As a result, the model develops stronger context understanding and instruction-following capabilities, which directly translate into improved controllability and perceptual quality in the voice design task.
@@ 3.5 Ablation Study of RAG
Table 6 demonstrates the necessity and effectiveness of incorporating RAG into the proposed VoiceSculpt-
or-VD framework. Enabling RAG leads to substantial improvements across all evaluation metrics, with especially pronounced gains in APS (+7.1%), RP (+13.0%), and the overall AVG score (+8.2%). These results indicate that external retrieval provides critical contextual and attribute-related information that the model alone may not reliably infer from the instruction text, thereby significantly enhancing controllability and instruction adherence in the voice design task.
@@
[…]
At the same time, the significant performance gap between the RAG and non-RAG settings also reveals inherent limitations of the current model. In particular, the model exhibits relatively limited text understanding and generalization when relying solely on its internal representations, making it sensitive to instruction phrasing and less robust to unseen or complex textual descriptions. By supplementing the model with retrieved examples and structured attribute information, RAG effectively compensates for these shortcomings.
4 Limitations and Future Work
The current model still exhibits several limitations in practical use. First, the model may exhibit limited stability, where repeated synthesis under the same instruction occasionally fails to maintain precise control over the desired attributes. Second, during speech synthesis and interaction, the system occasionally suffers from long periods of silence or delayed responses, which negatively affect the user experience. In terms of training data coverage, the model does not yet perform well for elderly and child voices, and the naturalness and timbre consistency for these groups remain insufficient.
Moreover, our current evaluation of the model’s instruction-following ability is incomplete. Existing benchmarks focus primarily on Chinese tasks, and we have not yet conducted a comprehensive assessment of their English or multilingual capabilities.
In future work, we plan to address these limitations more fundamentally by strengthening the model’s text understanding capabilities. Specifically, we will explore incorporating large-scale text data during the pre-training stage to preserve better and enhance linguistic representations, performing instruction data augmentation to improve robustness and generalization, and adopting more semantically expressive audio representations by replacing the current xcodec2. In addition, we will introduce larger-scale and more diverse data to further improve the model’s adaptability and expressiveness across different voice characteristics. These improvements are expected to reduce the model’s reliance on external retrieval while maintaining strong instruction-following performance.
@@ 5 Conclusion
In this work, we present VoiceSculptor, a unified natural-language-driven voice design framework that enables fine-grained, controllable speech synthesis. Extensive evaluations on the InstructTTSEval-Zh benchmark demonstrate that VoiceSculptor-VD achieves SOTA performance among open-source instruction-following TTS systems, consistently outperforming strong baselines across multiple complementary metrics. Our scaling study confirms that instruction-following capability benefits predictably from increased model capacity, richer training data, and staged training strategies with large-scale continual pre-training. Through systematic ablation studies, we further validate the effectiveness of key design choices, including CoT-based fine-grained attribute tokens, text-side cross-entropy supervision, and retrieval-augmented generation, each of which contributes to improved instruction understanding, controllability, and robustness. In particular, CoT-based attribute modeling enables more precise and interpretable control over vocal characteristics. At the same time, RAG effectively compensates for the model’s limited text generalization by providing semantically aligned in-domain guidance. Finally, we demonstrate that the voice design outputs of VoiceSculptor can be reliably transferred to downstream speech synthesis models, supporting practical deployment scenarios where voice design and speech generation are decoupled. Together, these results establish VoiceSculptor as a scalable and effective solution for natural-language-driven voice design.
@@ 6 Ethics Statement
Do not use this model for unauthorized voice cloning, impersonation, fraud, scams, deepfakes, or any illegal or malicious activities. Ensure compliance with local laws and regulations when using this model and uphold ethical standards. The developers assume no liability for any misuse of this model. Necessary clarification regarding generated voices: As a generative model, the voices produced by this system are synthetic outputs inferred by the model, not recordings of authentic human voices. The generated voice characteristics do not represent or reproduce any specific real individual, and are not derived from or intended to imitate identifiable persons. We advocate for the responsible development and use of AI and encourage the community to uphold safety and ethical principles in AI research and applications.